Problem 3-35 Using ratios to construct financial statements [LO3-2]

The following information is from Harrelson Inc.'s, financial statements. Sales (all credit) were $18.00 million for last year.    

Sales to total assets 1.20 times
Total debt to total assets 45 %
Current ratio 2.80 times
Inventory turnover 6 times
Average collection period 24 days
Fixed asset turnover 5 times

   
Complete the balance sheet: (Use a 360-day year. Do not round intermediate calculations. Input your answers in millions rounded to 2 decimal places.)
  

$ in millions $ in millions
Cash $1,000,000.00 Current debt
Accounts receivable 1,000,000.00 Long-term debt
Inventory Total debt
Total current assets
Fixed assets Equity
Total assets Total debt and stockholders' equity
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Sort &Find & 2 ClearFe Select Edting Format Painter Formatting, as Table w styles. ▼ ㆆ ▼ Clipboard Alignment Number Cells LG7 KY KZ LA LB LC LD LE LF LG LH LK Ll SALES 18 4 1 SALES TO TOTAL ASSETS SALES TOTAL ASSETS 1.2 = 18/TOTAL ASSETS 18/1.2 = 6 TOTAL ASSETS 15 8 2 TOTAL DEBT/TOTAL ASSETS = 45% TOTAL DEBT/ 15000000 = 45% TOTAL DEBT= 15 X 45% 10 6.75 12 13 14 15 16 17 18 19 3 INVENTORY TURNOVER= SALES/INVENTORY 18/INVENTORY INVENTORY- 4 FIXED ASSETS TURNOVER = SALES/ FIXED ASSETS 5 = 18/FIXED ASSETS FIXED ASSETS-18/5 = 3.6 | FIFO . CASH BUDGET 45 BV MV ratio VARIANCE BEP, OL FL ratios B-s t loss SALES BUDGET DIFF ANALYSIS . overhead . float erences: EV23 31-01-2019Home nert Page Layout Formulas Data Review View dd-Ins Cut E AutoSum Wrap Text ta copy ▼ в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Paste Sort &Find & Format Painter Formatting as Table Styles2 Clear Clipboard LG33 KY Alignment Number Cells Editing KZ LA LB LC LD LE LF LG LH ㄩ LK Ll 15 16 17 18 19 20 21 4 FIXED ASSETS TURNOVER = SALES/ FIXED ASSETS 5 18/FIXED ASSETS FIXED ASSETS-18/5 = 3.6 5 CURRENT ASSETS = TOTAL ASSETS-FIXED ASSETS CURRENT ASSETS = 15-3.6 CURRENT ASSETS = 11.4 23 24 25 26 27 28 29 30 31 32 6 AVERAGE COLLECTION PERIOD = 360 X ACCOUNT RECEIVABLE/SALES 24 = 360 X ACCOUNT RECEIVABLE/18 CCOUNT RECEIVABLE-18 X 24/360 - 7 CURRENT RATIO CURRENT ASSETS/CURRENT LIABILTIES 2.80 11400000/CURRENT LIABILITIES CURRENT LIABILITES 11.4/2.8- 4.07 BALANCE SHEET | FIFO . CASH BUDGET 45 BV MV rato VARIANCE BEP, OL FLratios B-S t loss SALES BUDGET DIFF ANALYSIS . overhead . float erences: EV23 05:33 31-01-2019Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text aCopy в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Paste Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard LG49 KY Alignment Number Cells Edting KZ LA LB LC LD LE LF LG LH LK Ll 31 32 BALANCE SHEET 34 35 36 37 38 39 40 41 42 43 CASH ACCOUNTS RECEIVABLE INVENTORY TOTAL CURRENT ASSETS 7.20 LC38-LC37-LC36 CURRENT DEBT 1.20 3.00 4.07 2.68 LG37-LG35 6.75 LONG TERM DEBT TOTAL DEBT 11.40 EQUITY 8.25 LG42-LG37 FIXED ASSETS 3.60 TOTAL ASSETS 15.00 15.00 TOTAL LIABILITIES & EQUITY 45 46 47 48 49 | FIFO . CASH BUDGET 45 BV MV rato VARIANCE BEP, OL FLratios B-S t loss SALES BUDGET DIFF ANALYSIS . overhead . float erences: EV23 05:33 31-01-2019